OpenStudy (anonymous):

simplify 1/(1+(3/2)^x)

OpenStudy (anonymous):

\[\frac{ 1 }{ 1+(3/2)^x }\]

OpenStudy (anonymous):

i did, \[\frac{ 1 }{ 1 } +\frac{ 1 }{ (3/2)^x } = 1+(\frac{ 2 }{ 3 })^x \], is that true?